A diligent business student, Bob, has decided to apply for an MBA program. The program says it will take a score from either the GRE or GMAT, but not both. Bob has taken both and is trying to decide which score is the better one to send. The information he has is below. Which score should Bob send? Justify your answer using a statistical argument. 4) GRE GMAT 695 Bobs Score Range of Scores 90th Percentile Mean (u) Standard Deviation (a) 160.5 Co 130-170 200-800 161.9 151.3 8.7 710 561 120
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

4. The score which is relatively higher w.r.t to own distribution should be sent.

Let' find that, Z score will help us out here.

Lets calculate Z scores for GRE and GMAT:

Zgre = (Bob' GRE score - Mean score)/Stdev = (160.5-151.3)/8.7 = 1.06

Zgmat = (Bob' Gmat score - Mean score)/Stdev = (695-561)/120 = 1.12

Since his Zgmat>Zgre, he has done relatively better on the GMAT than on the GRE.

So, BOB should send his GMAT score